I don't know how to show the steps for the equation. question: the sum of the digits of a two number is 15. If the digit in the tenth place is 8, find the two digit number. Can you help me understand the question .
Answer by fcabanski(1391)

A number is xy (like 15 where x is 1 and y is 5). x and y are the digits with y in the ones place and x in the tens place.


The problem indicates that the tens digit is 8, so the number is 8y where 8+y = 15.


Subtract 8 from both sides to find y=7.


The number is 87.


In the problem you indicated that 8 was in the "tenths" place. If it really is tenths place, and not tens place, then the number is x.y where y (the tenths place) is 8.


The other number is still x=15-8=7, but the number is 7.8.

I provide online tutoring ($30/hr) and personal problem solving
($3.50-$5.50 per problem. Contact me (check the profile "website" which is
my email address.